### v2

# Joint Beamforming and Position Optimization for Fluid RIS-aided ISAC Systems

Abstract
A fluid reconfigurable intelligent surface (fRIS)-aided integrated sensing and communications (ISAC) system is proposed to enhance multi-target sensing and multi-user communication. Unlike the conventional RIS, the fRIS incorporates movable elements whose positions can be flexibly adjusted to provide extra spatial degrees of freedom. In this system, a joint optimization problem is formulated to minimize sensing beampattern mismatch and communication symbol estimation error by optimizing the symbol estimator, transmit beamformer, fRIS phase shifts, and element positions. To solve this problem, an algorithm based on alternating minimization is devised, where subproblems are solved leveraging augmented Lagrangian method, quadratic programming, semidefinite-relaxation, and majorization-minimization techniques. A key challenge exists that the fRIS element positions affect both the incident and reflective channels, leading to the high-order composite functions regarding the positions. As a remedy, it is proved that the high-order terms can be transformed to linear and linear-difference forms using the characteristics of fRIS and structural channels, which facilitates the position optimization. Numerical results validate the effectiveness of the proposed scheme as compared to the conventional RIS-aided ISAC systems and other benchmarks.
Index Terms:
Alternating minimization, fluid reconfigurable intelligent surface, integrated sensing and communications, mismatch and estimation error
I Introduction
In the era of sixth-generation (6G) wireless communications, integrated sensing and communications (ISAC) emerged as a promising paradigm [1], addressing spectrum scarcity and reducing resource consumption. The ISAC aims to incorporate the sensing systems and the communication systems at different levels. Herein, the two systems shared the same spectrum and coexisted with minimal mutual interference [2] using null space projection [3], transceiver design [4], etc. To further mitigate interference, collaborations between the two systems were achieved by sharing critical information    [5], [6]. Additionally, considering the similarities in hardware and signal processing, the recent ISAC advancement focuses on designing unified platforms that can perform the dual functions [7]. For example, in [8], the transmit beamformer was designed by combining weighted radar waveforms and communication symbols. Besides, a wideband scenario about joint design was considered in [9], where mutual information was utilized as a unified metric to design the ISAC waveform.
Meanwhile, reconfigurable intelligent surface (RIS) garnered great attention in the 6G innovation. Specifically, the RIS consists of numerous reflective elements, each of which is able to induce phase shifts to the incident signals. This enables the RIS to dynamically adapt the channel propagation environment by designing the phase shifts [10]. Leveraging this characteristic, work [11] demonstrated significant improvement in communication sum-rate in RIS-aided multi-user scenarios. Apart from this, it was also shown that the RIS can suppress the undesired signals, such as eavesdropping  [12] and jamming [13]. Additionally, the RIS was applied in other diverse scenarios, including radar detection [14], near-field communications [15], and wireless power transfer and communications [16].
To harness the combined benefits of ISAC and RIS, extensive research was conducted on RIS-aided ISAC systems [17]. In [18], authors investigated an echo signal-to-noise ratio (SNR) maximization problem in a RIS-aided ISAC system, which serves a communication user and tracks a target through joint design of the transmit precoder and RIS beamformer. Building on this, [19] extended this framework to multi-user and multi-target scenarios under practical constraints. Furthermore, authors in [20] and [21] considered the energy efficiency maximization for the RIS-aided ISAC systems without compromising the communication and sensing performance. Additionally, researchers examined various other applications of the RIS-aided ISAC systems, such as simultaneously transmitting and reflecting surface [22], unsupervised learning [23], and extended reality [24].
While the RIS-aided ISAC systems offer significant enhancements for 6G communications, there remains room for the improvement in spatial degrees of freedom (DoFs) of the systems and the reduction in system complexity. Particularly, the antennas of the conventional arrays are fixed after assembly, which restricts the spatial diversity [25]. On the other hand, to achieve excellent performance, a massive number of elements are required, which increases manufacturing costs and exacerbates high-dimensional optimization complexity. To address these limitations, movable antenna, also known as fluid antenna, becomes a promising candidate. Enabled by the advancement of software-controllable fluidic metamaterials, the movable antenna can dynamically adjust its position in real time, offering additional spatial DoFs  [26], [27]. Moreover, the enhanced performance brought by the movable antenna allows comparable gains with fewer elements, significantly reducing the computational complexity caused by high-dimensional optimization.
The potential of the fluid antenna system was extensively studied. For instance, researchers demonstrated that it can promote the capacity improvement in point-to-point communication systems [25] and enhance downlink sum-rate in multi-user communication scenarios [28]. Besides, it was applied in ISAC systems to jointly optimize transceiver designs and antenna positions, enhancing communication and sensing performance [29], [30], [31]. Meanwhile, efforts to integrate the RIS with the fluid antenna system was made, where authors in [32] and [33] analyzed the outage probability bounds and asymptotic distributions for fluid antenna-aided RIS systems. Nevertheless, these works still rely on fixed RIS elements and do not fully exploit the spatial DoFs of RIS elements. By contrast, [34] introduced movable RIS elements with flexible positioning and studied the elimination of phase distribution offset in the system. However, this work was limited to a simple single-antenna and single-user communication setup, while did not consider the system optimization problem. This leaves room for further exploration in the potential of the movable RIS elements in ISAC systems.
Against the aforementioned background, we are motivated to study an ISAC system aided by a fluid RIS (fRIS). In particular, fRIS, named after the fluid antenna system, implies that the elements of RIS are moveable. In this system, beams are formed at the fRIS to serve the communication users and illuminate the targets, where the sensing beampattern mismatch and communication decoding error are jointly minimized. This problem is solved by an alternating minimization (AM)-based algorithm, which incorporates augmented Lagrangian method (ALM), quadratic programming (quadprog), semidefinite-relaxation (SDR), and majorization-minimization (MM) technique. The main contributions of this work are summarized as:
  • As the movable elements can provide additional spatial DoFs compared to the conventional RIS, we propose a novel fRIS-aided ISAC system to improve the communication and sensing performance. Specifically, the ISAC transmit signal, fRIS beampattern for target illumination, communication symbol decoding and fRIS-related channels are characterized, and a joint sensing beampattern mismatch and communication decoding error minimization problem is formulated.
  • To tackle the resultant optimization problem, we devise an algorithm based on AM. First, we reformulate the original high-order problem into a more manageable form. Then, the problem is divided into several subproblems and dedicated optimization approaches are established for each subproblem. Eventually, the convergence and the computational complexity of the proposed algorithm are analyzed.
  • In the subproblem of optimizing the fRIS element positions, the optimization process is challenging since the changing RIS elements simultaneously affects the incident channel, the reflective channels. To address this, we leverage the characteristics of the fRIS and the structural channels to reduce the fourth-order terms of the element position to the linear and linear difference terms, so that the objective function of the subproblem can be simplified. This facilitates the optimization process.
  • Numerical results are demonstrated to validate the effectiveness of the proposed model. The results reveal a significant performance gain achieved by employing the fRIS in the ISAC system and the proposed scheme outperforms the benchmarks. Moreover, by employing the fRIS, the system performance loss caused by the reduction of the element number can be complemented, which alleviates the curse of dimensionality in the RIS optimization.
This paper is organized as follows. Section II introduces the model architecture of the fRIS-aided ISAC system, while formulates the symbol error and beampattern mismatch minimization problem. In section III, the AM-based algorithm for the resultant problem and the analysis of the proposed algorithm are presented in details. Section IV demonstrates the numerical results. Finally, conclusions are drawn in Section V.
[…]
II System Model and Problem Formulation
II-A System Model
Figure 1: A scenario of a fRIS-aided ISAC system that performs multi-target sensing and multi-user communication.
As depicted in Fig.1, we consider an ISAC system aided by a fRIS. Specifically, the ISAC base station (BS), equipped with a $M$-antenna uniform linear array (ULA), aims to simultaneously serve $K$ single-antenna communication users and sense $T$ point targets. Without loss of generalization, it is considered that the direct sensing and communication paths are blocked so that the fRIS is deployed to establish the virtual links and sensing paths. The fRIS has $N$ movable elements, where all the elements can move freely within a planar region $\mathcal{A}$ with the shape of $A\times A$. For $n$-th fRIS element, its position is denoted by its Cartesian coordinate $\mathbf{p}_{n}=\left[p_{x,n},p_{y,n}\right]^{\mathrm{T}}\in\mathcal{A}^{2\times 1}$. As a result, the positions of all the elements can be given compactly as $\mathbf{p}=\left[\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{N}\right]\in\mathcal{A}^{2\times N}$.
In this system, we denote the data symbols for the users as $\mathbf{s}_{c}\in\mathbb{C}^{K\times 1}$, which is precoded by a beamformer $\mathbf{W}\in\mathbb{C}^{M\times K}$ in the BS. This yields the precoded digital signal as
$\mathbf{x}=\mathbf{W}\mathbf{s}_{c}\in\mathbb{C}^{M\times 1}.$ (1)
The symbol $\mathbf{s}_{c}$ is considered statistically independent and auto-correlated, namely $\mathbb{E}[\mathbf{s}_{c}\mathbf{s}_{c}^{\mathrm{H}}]=\mathbf{I}_{K\times K}$. After converting the radio frequency domain, the signal $\mathbf{x}$ is emitted from the BS and propagates through the channel $\mathbf{G}\in\mathbb{C}^{N\times M}$ to the fRIS. The phase of the incident signal in the fRIS is altered adaptively by an adjustable phase shift matrix $\mathbf{\Theta}=\mathrm{Diag}\left(\theta_{1},\theta_{2},\cdots,\theta_{N}\right)$, in which $\left\{\theta_{n},\forall n=1,\dots,N\right\}$ represents the phase shift coefficient for $n$-th RIS element. As a result, the reflective signal at the fRIS can be given as
$\mathbf{v}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}.$ (2)
Subsequently, the reflective signal at the fRIS ought to form multiple beams to illuminate the targets and propagate through the reflecting channel $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$ to serve the communication users.
II-A1 Sensing Model
As we know, the illumination power to the targets is direction related to the echo SNR, where more target information can obtained if the echo SNR is high. As a consequence, beams should be formed at the fRIS towards the directions of the targets to illuminate the targets. To this end, one can design the correlation matrix of the reflective signal $\mathbf{v}$ to manipulate the beampattern of the reflective signal. Specifically, the relationship between the beampattern and the correlation matrix of the reflective signal can be expressed as
$\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ $=\mathbf{a}^{\mathrm{H}}(\phi,\psi,\mathbf{p})\mathbf{R}_{s}\mathbf{a}(\phi,\psi,\mathbf{p}),$ (3)
in which $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}$ is the covariance matrix of signal $\mathbf{v}$. The symbol $\mathbf{a}(\phi,\psi,\mathbf{p})$ represents the steering vector of the fRIS, which is not only dependent on azimuth angle $\phi$ and elevate angle $\psi$, but also the positions of the movable elements $\mathbf{p}$. For conciseness, we leave the detailed expression of $\mathbf{a}(\phi,\psi,\mathbf{p})$ in the subsection II-A3.
For beam design, a common metric is the beampattern similarity, where the mismatch between the practical beampattern $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ and the ideal pattern $\mathcal{P}_{d}(\phi,\psi)$ are measured to evaluate the target sensing performance. Specifically, the beampattern mismatch can be written as
$\varepsilon_{r}=\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2},$ (4)
in which $\beta$ is a scaling factor, while $I_{a}$ and $I_{e}$ denote the grid number of the azimuth angles and elevate angles, respectively. When $\varepsilon_{r}$ is minimized, the practical beampattern aligns with the ideal beampattern, such that the reflective signal can effectively illuminate the targets.
II-A2 Communication Model
Apart from illuminating the targets, the reflective signal $\mathbf{v}$ also propagates to the users for communication through the channel $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$. The signal received at user $k$ can be given as
$y_{c,k}=\mathbf{h}_{rc,k}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+n_{c,k},~{}~{}\forall k=1,\cdots,K,$ (5)
where $n_{c,k}$ represents the additive Gaussian white noise at the receiver of user $k$ with zero mean and $\sigma_{0}^{2}$ variance. To estimate the transmitted symbols from the received signal, an estimation factor $\omega$ is introduced, i.e.,
$\hat{s}_{c,k}=\omega y_{c,k}.$ (6)
Given the estimated communication symbols, a direct performance indicator for evaluating the communication performance is the mean square error (MSE) between the estimated symbol and the transmitted symbol. It can be formulated as
$\varepsilon_{c}$ $=\mathbb{E}[\|\mathbf{s}_{c}-\hat{\mathbf{s}}_{c}\|^{2}]$ (7a)
=$\overset{(a)}{=}\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2},$ (7b)
where $\mathbf{H}_{rc}=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,k}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}\in\mathbb{C}^{N\times K}$ incorporates all the reflective channels. The operation $(a)$ substitutes (5) and (6) to (7a) and leverages the statistical independence and the auto-correlation characteristics of $\mathbf{s}_{c}$.
II-A3 Channel Model
As stated before, the steering vector of the fRIS is dependent on the positions of the elements $\mathbf{p}$ as well as the azimuth angle $\phi$ and the elevate angle $\psi$. Specifically, the steering vector of the fRIS can be expressed as [25]:
$\!\!\mathbf{a}(\phi,\psi,\mathbf{p})\!=\!\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}d_{1}(\phi,\psi,\mathbf{p}_{1})}&\!\!\!\cdots\!\!\!&e^{\jmath\frac{2\pi}{\lambda}d_{N}(\phi,\psi,\mathbf{p}_{N})}\end{bmatrix}^{\mathrm{T}}\!\!\!.$ (8)
The symbol $d_{n}(\phi,\psi,\mathbf{p})$ is referred to as the path difference of the signal propagation from the direction $(\phi,\psi)$ between the $n$-th element and the reference origin, which gives
$d_{n}(\phi,\psi,\mathbf{p}_{n})=p_{x,n}\sin(\phi)\cos(\psi)+p_{y,n}\sin(\psi).$ (9)
Without loss of generalization, the fRIS is typically deployed in the area where line-of-sight (LoS) paths dominate, so that the signal coverage can be extended. Consequently, the fRIS-related channels are modeled as LoS channels, which are determined by the steering vector of the fRIS. For the channel $\mathbf{G}$ from the BS to the fRIS, it can be modeled as
$\mathbf{G}$ $=\sqrt{\zeta_{G}}\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}(\phi_{t}),$ (10)
in which $\sqrt{\zeta_{G}}$ denotes the path loss between the BS and the fRIS, while $\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})$ represents the receive steering vector of the fRIS with the signal coming from $\phi_{r}$ in azimuth direction and $\psi_{r}$ in elevate direction, which is in the form of (8). On the other hand, the ULA is equipped in the BS, so that the transmit steering vector $\mathbf{a}_{t}$ in the direction of $\phi_{t}$ can be written as
$\mathbf{a}_{t}=\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{1}}&\cdots&e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{m}}&\cdots&e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{M-1}}\end{bmatrix}^{\mathrm{T}},$ (11)
where $\{\widetilde{d}_{m}=\frac{\lambda}{2}m\sin(\phi_{t}),~{}\forall m=0,\cdots,M-1\}$ and the inter-element spacing is half wavelength. Similarly, the channel $\mathbf{h}_{rc,k}^{\mathrm{H}}$ from the fRIS to the user $k$ is modeled as
$\mathbf{h}_{rc,k}^{\mathrm{H}}=\sqrt{\zeta_{k}}\mathbf{a}^{\mathrm{H}}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p}),\forall k=1,\cdots,K,$ (12)
where $\mathbf{a}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p})$ denotes the transmit steering vector of the fRIS to user $k$ from azimuth direction $\phi_{c,k}$ and elevate direction $\psi_{c,k}$, which takes the form of (8). The path loss between the fRIS and the user $k$ is denoted as $\sqrt{\zeta_{k}}$. Accordingly, the overall reflective channel $\mathbf{H}_{rc}$ can be rewritten as
$\mathbf{H}_{rc}$ $=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}=\mathbf{A}_{rc}\boldsymbol{\Sigma}_{rc},$ (13)
where we denote
$\mathbf{A}_{rc}$ $=\begin{bmatrix}\mathbf{a}_{c,1}(\phi_{c,1},\psi_{c,1},\mathbf{p})&\!\!\!\cdots\!\!\!&\mathbf{a}_{c,K}(\phi_{c,K},\psi_{c,K},\mathbf{p})\end{bmatrix},$ (14a)
$\boldsymbol{\Sigma}_{rc}$ $=\mathrm{Diag}\left(\begin{bmatrix}\sqrt{\zeta_{1}}&\cdots&\sqrt{\zeta_{K}}\end{bmatrix}\right).$ (14b)
II-B Problem Formulation
In order to simultaneously achieve the sensing and the communication functions, a weighted sum of the sensing beampattern mismatch $\varepsilon_{r}$ of (4) and the communication symbol estimation MSE $\varepsilon_{c}$ of (7) is minimized by optimizing the transmit beamformer $\mathbf{W}$, the fRIS phase shift $\boldsymbol{\Theta}$, the fRIS element positions $\mathbf{p}$, the communication symbol estimator $\omega$ as well as the scaling factor $\beta$ of the sensing beampattern. Consequently, the problem can be formulated as
$\min_{\mathbf{W},\boldsymbol{\Theta},\mathbf{p},\omega,\beta}$ $~{}\varepsilon_{0}\!=\!\alpha\left(\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2}\right)$
$+(1\!-\!\alpha)\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!K\omega^{2}\sigma_{0}^{2}\right)$ (15a)
$\mathrm{s.t.}$ $~{}\|\mathbf{W}\|^{2}\leq P_{t},$ (15b)
$~{}\omega\in\mathbb{R},$ (15c)
$~{}|\boldsymbol{\Theta}|_{n,n}=1,\forall n=1,\cdots,N$ (15d)
$~{}\mathbf{p}_{n}\in\mathcal{A},\forall n=1,\cdots,N$ (15e)
$~{}\|\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}\!\geq\!\Delta D,\forall n,n^{\prime}\!\!=\!\!1,\!\cdots\!,N,n\!\neq\!n^{\prime},$ (15f)
where $\alpha$ is the weighting coefficient indicating the bias of the sensing and the communication. The constraint (15b) represents that the transmit power should not exceed $P_{t}$, while in the constraint (15d), the modulus-one phase shift of the fRIS is required. The constraints (15e) and (15f) indicate that each RIS element ought to move within the pre-set planar region, and the inter-space among elements should be larger than $\Delta D$. It is challenging to solve the problem (15) owing to the two reasons. On the one hand, (15d) is a constant modulus constraint, while (15f) takes the non-convex form of “Norm $\geq$ Constant”. Besides, the objective function is highly complex, where the term of beampattern mismatch $\varepsilon_{r}$ in the objective function is high-order with respect to different variables.
III Proposed Algorithm
In this section, an effective AM-based algorithm is developed to address the problem (15) by iteratively updating the variables, and we analyze the convergence and computational complexity of the proposed algorithm.
III-A Reformulation of $\varepsilon_{r}$
Carefully inspecting $\varepsilon_{r}$ in (4), it can be observed that $\varepsilon_{r}$ is a quadratic function with respect to the covariance matrix $\mathbf{R}_{s}$, while $\mathbf{R}_{s}$ is also a quadratic function with respect to the variables, making $\varepsilon_{r}$ a quartic function with respect to the variables. To simplify the objective function in (15a), we reduce the order of $\varepsilon_{r}$ to be a quadratic function with respect to the variables.
Specifically, we first design an optimum covariance matrix $\widetilde{\mathbf{R}}_{s}$ that only takes the sensing into consideration. According to [35], the problem of designing $\widetilde{\mathbf{R}}_{s}$ can given as
$\min_{\beta,\widetilde{\mathbf{R}}_{s}}$ $~{}\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}},\widetilde{\mathbf{R}}_{s})\right|^{2}$ (16a)
$\mathrm{s.t.}$ $~{}\beta>0,\widetilde{\mathbf{R}}_{s}\succeq 0,$ (16b)
$\mathrm{tr}\left(\widetilde{\mathbf{R}}_{s}\right)={\|\mathbf{G}\mathbf{x}\|^{2}},$ (16c)
$\mathrm{rank}\left(\widetilde{\mathbf{R}}_{s}\right)=1,$ (16d)
where the objective function requires the designed sensing beampattern to be similar to an ideal one, the constraint (16b) indicates that $\widetilde{\mathbf{R}}_{s}$ is semi-definite, the radiation power at the fRIS is $\|\mathbf{Gx}\|^{2}$ in the constraint (16c), and the last constraint requires $\widetilde{\mathbf{R}}_{s}$ to be rank-one since the ultimately designed $\mathbf{R}_{s}$ should satisfy $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}$. The problem (16) can be directly solved with SDR approach. To guarantee the rank-one constraint of $\widetilde{\mathbf{R}}_{s}$, various methods can be employed, such as Gaussian randomization [36] and penalty-based method [37].
With the obtained $\widetilde{\mathbf{R}}_{s}$ designed in (16), the beams can be effectively formed towards the targets. Considering that the designed $\widetilde{\mathbf{R}}_{s}$ is rank-one, the corresponding reflective signal at the fRIS associated with $\widetilde{\mathbf{R}}_{s}$ can be retrieved by the performing eigenvalue decomposition to $\widetilde{\mathbf{R}}_{s}$. This yields
$\widetilde{\mathbf{R}}_{s}=\mathbf{s}_{r}\mathbf{s}_{r}^{\mathrm{H}},$ (17)
where $\mathbf{s}_{r}=\sqrt{\lambda_{s}}\mathbf{u}_{s}$, while $\lambda_{s}$ and $\mathbf{u}_{s}$ represent the maximum eigenvalue and the corresponding eigenvector of $\widetilde{\mathbf{R}}_{s}$, respectively. The factorized signal $\mathbf{s}_{r}$ represents an optimum and achievable reflective signal whose correlation matrix is $\widetilde{\mathbf{R}}_{s}$ producing directive beams towards the targets. In other words, $\mathbf{s}_{r}$ can be viewed as a desired reflective signal for sensing the targets.
Consequently, the sensing metric is reformulated as MSE between the practical reflective signal $\mathbf{v}$ at the fRIS and the desired reflective signal $\mathbf{s}_{r}$, i.e.
$\widetilde{\varepsilon}_{r}=\|\mathbf{s}_{r}-\mathbf{v}\|^{2}=\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}.$ (18)
It can be observed that the reformulated sensing metric $\widetilde{\varepsilon}_{r}$ becomes quadratic with respect to all the variables, which simplifies the subsequent optimization.
Ultimately, the problem (15) can be reformulated as
$\min_{\mathbf{W},\omega,\mathbf{p},\boldsymbol{\Theta}}$ $~{}\varepsilon=\alpha\left(\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\right)$
$+(1\!-\!\alpha)\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!K\omega^{2}\sigma_{0}^{2}\right)$ (19a)
[…]
However, the problem is still challenging owing to the non-convex constraints and objective function. In the subsequence, the problem (19) is solved by iteratively updating each variable with others being fixed. It is worth mentioning that $\mathbf{s}_{r}$ may change after the update of $\mathbf{p}$, but according to the Lyapunov stability [38], $\mathbf{s}_{r}$ is asymptotically stable over iterations due to the bounded $\mathbf{p}$ and the finite resources.
III-B Optimization of Communication Symbol Estimator $\omega$
When $\mathbf{x}$, $\mathbf{p}$ and $\boldsymbol{\Theta}$ are fixed, the objective function in the problem (19) is a quadratic function with respect to $\omega$. After ignoring the terms that are unrelated to $\omega$, the problem can be simplified as
$\min_{\omega\in\mathbb{R}}$ $~{}~{}\varepsilon_{c}=\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2}.$ (20)
The optimal $\omega^{\star}$ can therefore be derived by letting the derivative of $\varepsilon_{c}$ be zero, that is, $\frac{\partial\varepsilon_{c}}{\partial\omega}=0$. This yields a closed-form optimum:
$\omega^{\star}=\frac{\mathfrak{R}\{\mathbf{s}^{\mathrm{H}}_{c}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}}{\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+K\sigma_{0}^{2}}.$ (21)
III-C Optimization of fRIS Phase Shift $\boldsymbol{\Theta}$
Given $\mathbf{W}$, $\mathbf{p}$ and $\omega$, the problem (19) can be reduced to a constant-modulus constrained quadratic minimization problem, which is given as
$\min_{\boldsymbol{\Theta}}$ $~{}\alpha\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!(1\!-\!\alpha)\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}.$ (22a)
$\mathrm{s.t.}$ $~{}|\boldsymbol{\Theta}|_{n,n}=1,~{}~{}\forall~{}n=1,\cdots,N.$ (22b)
The challenge of this problem mainly stems from the non-convex constant-modulus constraint. To address this task, one can first let $\boldsymbol{\theta}=\mathrm{diag}(\boldsymbol{\Theta})$ and employ the property of $\mathrm{tr}(\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{C}_{1}\boldsymbol{\Theta}\mathbf{C}_{2})=\boldsymbol{\theta}^{\mathrm{H}}(\mathbf{C}_{1}\odot\mathbf{C}_{2}^{\mathrm{T}})\boldsymbol{\theta}$ [21] to transform the problem (22) to a compact form:
$\min_{\boldsymbol{\theta}}$ $~{}\boldsymbol{\theta}^{\mathrm{H}}{\mathbf{A}}_{1}\boldsymbol{\theta}-\boldsymbol{\theta}^{\mathrm{H}}{\mathbf{b}}_{1}-{\mathbf{b}}_{1}^{\mathrm{H}}\boldsymbol{\theta}$ (23a)
$\mathrm{s.t.}$ $~{}|\boldsymbol{\theta}|_{n}=1,\forall n=1,\cdots,N.$ (23b)
where
${\mathbf{A}}_{1}$ $\!=\!\alpha\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\odot\mathbf{I}\right]$
$~{}~{}+\omega^{2}\left(1\!-\!\alpha\right)\!\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\!\odot\!\left(\mathbf{H}_{rc}^{*}\mathbf{H}_{rc}^{\mathrm{T}}\right)\right],$ (24a)
${\mathbf{b}}_{1}$ $=\alpha\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}})+\omega(1-\alpha)\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}).$ (24b)
The nonconvex constant-modulus constraint (23b) remains challenging, where various approaches have been developed in the wide literature to effectively handle such a constraint, such as SDR [39] and manifold optimization [40]. Therefore, the details will not be elaborated upon here.
III-D Optimization of Transmit Beamformer $\mathbf{W}$
In this subsection, the optimization of the transmit beamformer $\mathbf{W}$ will be discussed. Viewing other variables as constant, the problem (19) can be reduced to
$\min_{\mathbf{W}}$ $~{}\alpha\!\left(\!\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}\|^{2}\right)\!\!+\!(1\!-\!\alpha)\!\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}\|^{2}\right)$ (25a)
$\mathrm{s.t.}$ $~{}\|\mathbf{W}\|^{2}\leq P_{t},$ (25b)
where we denote $\mathbf{H}_{c}=\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ as the overall channel from the transmitter to the users. For the objective function (25a), the norm operations are expanded and thus it becomes
$\alpha\!\left(\mathbf{s}_{c}^{\mathrm{H}}\mathbf{W}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}\}\right)$
$+$ $(1\!-\!\alpha)\!\left(\omega^{2}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{W}^{\mathrm{H}}\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}-2\omega\mathfrak{R}\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}\}\right).$ (26)
By leveraging the cyclic property of trace, the relationship of $\mathrm{tr}(\mathbf{X}^{\mathrm{H}}\mathbf{C}_{1}\mathbf{X}\mathbf{C}_{2})=\mathrm{vec}^{\mathrm{H}}(\mathbf{X})(\mathbf{C}_{2}^{\mathrm{T}}\otimes\mathbf{C}_{1})\mathrm{vec}(\mathbf{X})$ and $\mathrm{tr}(\mathbf{C}^{\mathrm{H}}\mathbf{X})=\mathrm{vec}^{\mathrm{H}}(\mathbf{C})\mathrm{vec}(\mathbf{X})$, the objective function can further be simplified as
$\mathbf{w}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{w}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{w}\},$ (27)
in which
$\mathbf{w}$ $=\mathrm{vec}(\mathbf{W}),$ (28a)
$\mathbf{A}_{2}$ $\!=\!\alpha\!\left(\mathbf{s}_{c}^{*}\mathbf{s}_{c}^{\mathrm{T}}\!\otimes\!\mathbf{G}^{\mathrm{H}}\mathbf{G}\right)\!+\!(1\!-\!\alpha)\omega^{2}\!\left(\mathbf{s}_{c}^{*}\mathbf{s}_{c}^{\mathrm{T}}\!\otimes\!\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\right),$ (28b)
$\mathbf{b}_{2}$ $=\alpha\left(\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{s}_{r}\mathbf{s}_{c}^{\mathrm{H}}\right)+(1\!-\!\alpha)\omega\left(\mathbf{H}_{c}^{\mathrm{H}}\mathbf{s}_{c}\mathbf{s}_{c}^{\mathrm{H}}\right).$ (28c)
It is worth noting that in (28b), $\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ has been simplified as $\mathbf{G}^{\mathrm{H}}\mathbf{G}$ since $\boldsymbol{\Theta}$ is a modulus-one diagonal matrix and $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ always holds. After this transformation, it can be seen that the objective (27) is a convex quadratic function. As for the constraint (25b), it can also be re-expressed as the inner product of $\mathbf{w}$, that is, $\mathbf{w}^{\mathrm{H}}\mathbf{w}\leq P_{t}$. In this context, we consider full power usage in the transmitter, namely $\mathbf{w}^{\mathrm{H}}\mathbf{w}=P_{t}$. As a result, the problem (25) can be written as
$\min_{\mathbf{w}}$ $~{}\mathbf{w}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{w}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{w}\}$ (29a)
$\mathrm{s.t.}$ $~{}\mathbf{w}^{\mathrm{H}}\mathbf{w}=P_{t},$ (29b)
which can be addressed by ALM [41, 42]. Specifically, we first employ the complex-real conversion by separating the real part and the imaginary part. The problem can thus be converted to its real form as
$\min_{\overline{\mathbf{w}}}$ $~{}\mathring{f}(\overline{\mathbf{w}})=\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{A}}_{2}\overline{\mathbf{w}}-2\overline{\mathbf{b}}_{2}^{\mathrm{T}}\overline{\mathbf{w}}$ (30a)
$\mathrm{s.t.}$ $~{}\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}=P_{t},$ (30b)
where
$\overline{\mathbf{w}}$ =[$=\begin{bmatrix}\mathfrak{R}\{\mathbf{w}\}\\ \mathfrak{I}\{\mathbf{w}\}\end{bmatrix},\overline{\mathbf{b}}_{2}=\begin{bmatrix}\mathfrak{R}\{\mathbf{b}_{2}\}\\ \mathfrak{I}\{\mathbf{b}_{2}\}\end{bmatrix},$ (31a)
$\overline{\mathbf{A}}_{2}$ =[$=\begin{bmatrix}\mathfrak{R}\{\mathbf{A}_{2}\}&-\mathfrak{I}\{\mathbf{A}_{2}\}\\ \mathfrak{I}\{\mathbf{A}_{2}\}&\mathfrak{R}\{\mathbf{A}_{2}\}\end{bmatrix}.$ (31b)
Accordingly, the augmented Lagrangian function of problem (30) can be written as
$\mathcal{L}\triangleq\mathring{f}(\overline{\mathbf{w}})+\mu\left(\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right)+\frac{\gamma}{2}\left(\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right)^{2},$ (32)
in which $\mu$ is the Lagrange multiplier and $\gamma$ denotes the penalty factor. This can be tackled by the following iteration procedure:
$\overline{\mathbf{w}}^{(t)}$ $=\arg\min_{\overline{\mathbf{w}}}\mathcal{L}(\overline{\mathbf{w}},\mu^{(t-1)}),$ (33a)
$\mu^{(t)}$ $=\mu^{(t-1)}+\gamma\left(\overline{\mathbf{w}}^{\mathrm{T}(t)}\overline{\mathbf{w}}^{(t)}-P_{t}\right).$ (33b)
Specifically, one can solve the subproblem (33a) by quasi-Newton method [42], which is implemented by the fminunc tool. This process is terminated when $\left|\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right|<\epsilon_{0}$ is satisfied. It is noted that the penalty factor $\gamma$ gradually increases among the iterations for fast convergence. After obtaining $\overline{\mathbf{w}}^{\star}$, it is converted to the complex matrix form and reshaped as $\mathbf{W}^{\star}$.
III-E Optimization of fRIS Element Positions $\mathbf{p}$
In this subsection, we focus on the optimization of the fRIS element position $\mathbf{p}$ with other variables being fixed. First, the problem (19) can be simplified by omitting the unrelated terms and constraints:
$\min_{\mathbf{p}}$ $~{}\alpha\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}(\mathbf{p})\mathbf{x}\|^{2}$
$~{}+(1-\alpha)\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}(\mathbf{p})\mathbf{x}\|^{2}$ (34a)
$\mathrm{s.t.}$ $~{}\mathbf{p}_{n}\in\mathcal{A},\forall n=1,\cdots,N,$ (34b)
$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}\geq\Delta D,\forall n,n^{\prime}=1,\cdots,N,n\neq n^{\prime}.$ (34c)
As mentioned in section II-A3, the channel $\mathbf{G}$ and $\mathbf{H}_{rc}$ are determined by the fRIS element positions. It is challenging to solve the problem (34) owing to two main reasons. First, the objective function is a complicated composite function concerning the variable $\mathbf{p}$. Since the fRIS element positions affect both the incident channel $\mathbf{G}$ and the reflective channel $\mathbf{H}_{rc}$, there appear quadratic and quartic forms simultaneously in the outer function. In the interior function, the variable $\mathbf{p}$ is contained in the complex exponential function of an array manifold, shown in (8). On the other hand, the constraint (34c) takes the form of “Norm $\geq$ Constant”, rendering it non-convex.
To explicitly observe $\mathbf{p}$ in (34a), we process the outer function by substituting $\mathbf{G}$ in (10) and $\mathbf{H}_{rc}$ in (11) to (34a) and integrating the unrelated symbols of $\mathbf{p}$, which is given by the following proposition.
Proposition 1. After omitting the constants, the objective function (34a) can be rewritten as
$\widetilde{f}_{0}=$ $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\right\}$
$+(1\!-\!\alpha)\left[c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\right.$
$\left.-2\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathrm{tr}\left(\mathbf{S}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{r}(\mathbf{p})\right)\right\}\right],$ (35)
where we denote $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$, $c_{0}=\omega^{2}\zeta_{G}c_{x}$, $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$, $\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{c,k}(\mathbf{p}))$, $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
Proof: Detailed derivations are provided in Appendix A.
However, the variable in (III-E) is contained in the array manifold, while there still exist $K$ quartic function terms, a quadratic function term and a linear function item with respect to the array manifold. Fortunately, it can be proved in the following proposition that the quartic, quadratic and linear items in (III-E) can be reduced to the first-order items and the difference of the first-order items if we only focus on optimizing the position of one element at a time.
Proposition 2. When considering the position of $n$-th element $\mathbf{p}_{n}$ while fixing others, the objective function $\widetilde{f}_{0}$ in (III-E) can be further simplified as
$\widetilde{f}_{1}(\mathbf{p}_{n})$ $=\widetilde{f}_{1,1}+\widetilde{f}_{1,2}+\widetilde{f}_{1,3}$
$=-\nu_{n}\cos\left(\xi_{n}+\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})\right)$
$~{}~{}+\sum_{k=1}^{K}\sum_{i=1,i\neq n}^{N}\widetilde{\nu}_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right)$
$~{}~{}-\sum_{k=1}^{K}\overline{\nu}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right).$ (36)
In the first term, we denote $\xi_{n}$ and $\rho_{n}$ as the phase and amplitude of $n$-th entry of ${}\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\nu_{n}\!=\!2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is the constant. The symbol $d_{r,n}(\mathbf{p}_{n})$ is $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})\!+\!p_{y,n}\sin(\psi_{r})$. In the second term, $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\!=\!d_{r,n}(\mathbf{p}_{n})\!-\!d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\xi}_{i,n,k}$ is the phase of a constant $\widetilde{c}_{i,n,k}$. Meanwhile, we denote $\widetilde{\nu}_{k}=2(1\!-\!\alpha)c_{0}\zeta_{k}$. In the last term, we represent $\overline{\xi}_{n,k}$ as the phase of $s_{n,k}$ and $\overline{\nu}_{n,k}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\overline{\rho}_{n,k}$.
Proof: Detailed derivations and symbol explanations are shown in Appendix B.
Observing the objective function $\widetilde{f}_{1}$, it contains cosine functions, which are neither convex nor concave with respect to $\mathbf{p}_{n}$. To address this problem, we employ the MM framework, where surrogate functions that locally upper bound the objective function $\widetilde{f}_{1}$ are constructed. Specifically, the constructions of the surrogate function is given by the following lemma.
Lemma 1. [43] Considering a function $\widetilde{f}$ with $\mathbf{x}$ being the variable, its second-order Taylor expansion at the given point $\mathbf{x}_{0}$ can be given as
$\widetilde{f}(\mathbf{x})\approx\widetilde{f}(\mathbf{x}_{0})$ $+(\nabla\widetilde{f}(\mathbf{x}_{0}))^{\mathrm{T}}(\mathbf{x}-\mathbf{x}_{0})$
$+\frac{1}{2}(\mathbf{x}-\mathbf{x}_{0})^{\mathrm{T}}\nabla^{2}\widetilde{f}(\mathbf{x}_{0})(\mathbf{x}-\mathbf{x}_{0}),$ (37)
where $\nabla\widetilde{f}(\mathbf{x})$ is the gradient vector of $\widetilde{f}(\mathbf{x})$, while $\nabla^{2}\widetilde{f}(\mathbf{x}_{0})$ represents the Hessian matrix of the function.
By selecting a parameter $\delta$ satisfying $\delta\mathbf{I}\succeq\nabla^{2}\widetilde{f}(\mathbf{x}_{0})$, the function $\widetilde{f}(\mathbf{x})$ can be locally upper-bounded by
$\!\widetilde{f}(\mathbf{x})\leq\widetilde{f}(\mathbf{x}_{0})\!+\!(\nabla\widetilde{f}(\mathbf{x}_{0}))^{\mathrm{T}}(\mathbf{x}-\mathbf{x}_{0})\!+\!\frac{\delta}{2}\left\|(\mathbf{x}-\mathbf{x}_{0})\right\|^{2}.$ (38)
In particular, the parameter $\delta$ can be chosen as the Frobenius norm of the Hessian matrix $\nabla^{2}f(\mathbf{x}_{0})$.
According to the Lemma 1, we separate $\mathbf{p}_{n}$ from the cosine functions and utilize the second-order Taylor expansion as the surrogate function. To construct the surrogate function, the gradient vector and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ should be determined. Specifically, the gradient and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ are given by the following theorem.
Theorem 1. According to the linear property of derivative operation, the gradient of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the derivatives of its three components, which can be expressed as
$\nabla\widetilde{f}_{1}(\mathbf{p}_{n})=\begin{bmatrix}\sum_{i=1}^{3}\frac{\partial\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{x,n}}&\sum_{i=1}^{3}\frac{\partial\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{y,n}}\end{bmatrix}^{\mathrm{T}}.$ (39)
For the Hessian matrix of $\widetilde{f}_{1}$, it can also be given as
$\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})=\begin{bmatrix}\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p^{2}_{x,n}}&\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{x,n}\partial p_{y,n}}\\ \sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{y,n}\partial p_{x,n}}&\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p^{2}_{y,n}}\end{bmatrix}.$ (40)
Specifically, there are 6 first-order derivative terms in (39) and 12 second-order derivative terms in (40), whose detailed expressions are left in the Appendix C for brevity.
Proof: Please refer to Appendix C.
According to Lemma 1 and Theorem 1, the equivalent objective function ultimately be given as
$(\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)}))^{\mathrm{T}}(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\!+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\right\|^{2},$ (41)
where the constants are neglected and $\delta_{n}=\|\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})\|^{2}_{\mathrm{F}}$. The $\mathbf{p}_{n}^{(t-1)}$ represents the output position of $n$-th element from the last round optimization. Apparently, the function (41) is a convex quadratic function with respect to $\mathbf{p}_{n}$.
Apart from the objective function, the constraint (34c) remains non-convex in the current form. To deal with it, an effective approach is to relax the constraint (34c) with the first-order Taylor expansion. Specifically, the first-order Taylor expansion of the left term in (34c) can be given as
$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}$ $\geq\|\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}$
$~{}~{}~{}+\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (42a)
$=\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (42b)
As a result, the constraint (34c) can be relaxed as
$\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}\!\geq\!\Delta D,\forall n,n^{\prime}\!=\!1:N,n\!\neq\!n^{\prime},$ (43)
which is affine regarding $\mathbf{p}_{n}$. Therefore, the problem for optimizing the position of the $n$-th element is given as
$\min_{\mathbf{p}_{n}}$ $~{}(\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)}))^{\mathrm{T}}(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\right\|^{2}$ (44a)
$\mathrm{s.t.}$ $~{}\mathbf{p}_{n}\in\mathcal{A},$ (44b)
$\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}\!\geq\!\Delta D,\forall n\!\neq\!n^{\prime}.$ (44c)
The problem (44) is a typical convex quadratic problem, which can be solved by using quadprog [44, 45]. Besides, we adopt circle packing scheme [46] for the initialization of fRIS element positions, where the initial positions are sufficiently separated. The details of the proposed algorithm are summarized in Algorithm 1.
Algorithm 1 Optimization of fRIS element positions
0:    Channel parameters, transmit signal, fRIS phase shift matrix, system parameters, desired reflective signal and communication symbols;
1:  Initialize the element positions $\mathbf{p}^{(0)}$ with the circle packing scheme;
2:  for $n=1\longrightarrow N$ do
3:     Calculate the gradient and the Hessian matrix of $\widetilde{f}_{1}$ according to the Theorem 1;
4:     Calculate $\delta_{n}$ based on the Hessian matrix of $\widetilde{f}_{1}$;
5:     Solve the problem (44) using quadprog;
6:  end for
7:  return  fRIS element positions $\mathbf{p}$.
III-F Convergence and Computational Complexity Analysis
The proposed algorithm is summarized in Algorithm 2, which incorporates the solutions derived above. The algorithm iteratively optimizes the variables, where the process continues until the objective function value in (19) becomes stable.
III-F1 Convergence Analysis
In this part, we analyze that the convergence of the proposed Algorithm 2. First, the subproblem (20) is a unconstrained convex quadratic problem with a close-form solution, and thus $\omega^{t}$ update does not increase objective function value:
$\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t}).$ (45)
Updating $\mathbf{\Theta}^{t}$ in (23), no matter utilizing the SDR [39] or the manifold method [40], it was proved that the objective function value of (23) is non-increasing over iterations, such that
$\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t+1},\mathbf{W}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t}).$ (46)
In the optimization of $\mathbf{W}$, quasi-Newton method guarantees that the update of $\overline{\mathbf{w}}$ is non-increasing [42], while equality constraint is satisfied if the penalty parameter becomes sufficiently large. As a result, the solution to (33) can satisfy the equality constraint, while the process (33) is non-increasing:
$\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t}\!,\mathbf{p}^{t}).$ (47)
Proved in [25], the subproblem (44) is a non-decreasing process given the nature of its convexity and the inequalities of Taylor expansions, while the constrained optimization process can converge in the MM framework when the function is smooth objective function [47]:
$\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t+1})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t}).$ (48)
Besides, the finite resources indicate that the objective function value is bounded and greater than zero. As a result, the Algorithm 2 can guarantee to converge.
III-F2 Computational Complexity Analysis
In this part, the computational complexity is analyzed.
  • To obtain $\mathbf{s}_{r}$, the computational complexity mainly stems from solving the problem (16) with SDR and the eigenvalue decomposition. Given a solution accuracy $\epsilon_{1}$, the computational complexity of SDR is $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$ [36], while the eigenvalue decomposition has the computational complexity of $\mathcal{O}\left(N^{3}\right)$.
  • The optimal symbol estimator can be calculated via a close-form expression, whose computational complexity stems from matrix multiplications. Obtaining the numerator in (21) has the computation complexity of $\mathcal{O}\left(KN+N^{2}+MN+M\right)$, while calculating the denominator has the computation complexity of $\mathcal{O}\left(2(MK+NK+N^{2})+M\right)$.
  • When optimize the fRIS phase shift, the computational complexity of calculation of the matrix multiplication is $\mathcal{O}\left(MN+MK+N^{2}K+N^{2}M+MNK\right)$. Besides, the SDR approach has the computational complexity of $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{2})\right)$ with $\epsilon_{2}$ being the solution accuracy, while that of the manifold optimization is $\mathcal{O}\left(N^{1.5}\right)$ [40].
  • In the optimization of transmit beamformer design, the computational complexity of matrix multiplication is $\mathcal{O}\left(M^{2}K^{2}\!+\!M^{2}N\!+\!M^{2}K\!+\!N^{2}M\!+\!MK\!+\!MN\!+\!K^{2}\right)$, while the design of the transmit beamformer with ALM is $\mathcal{O}\left((M^{2}+M)\log\left(1/\epsilon_{3}\right)\right)$ with $\epsilon_{3}$ being the convergence accuracy [48, 49].
  • When optimize the position of one fRIS element, the computational complexity of obtaining the gradients and the Hessian matrix is $\mathcal{O}\left(NK+K+M+N^{2}\right)$, while the computational complexity of calculating $\delta_{n}$ is $\mathcal{O}\left(1\right)$. Additionally, the computational complexity of solving the problem (44) is $\mathcal{O}\left(\left(N^{1.5}\log(1/\epsilon_{4})\right)\right)$ [25] with accuracy $\epsilon_{4}$ for the interior-point method.
As a result, the total computational complexity of each iteration of the Algorithm 2 can be approximated as $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$, neglecting the lower-order terms.
Algorithm 2 Proposed alternating optimization algorithm
0:    Channel parameters, noise power, transmit power budget, ideal sensing beampattern, weighting coefficient, users number, fRIS planar region size, and the minimum inter-element spacing requirement.
1:  Initialize all the variables;
2:  while no convergence of $\varepsilon$ do
3:     Calculate correlation matrix for sensing-only by (16) and obtain desired reflective signal $\mathbf{s}_{r}$ via (17);
4:     Update communication symbol estimator by (21);
5:     Optimize $\boldsymbol{\Theta}$ by solving problem (23);
6:     Update ${\mathbf{W}}$ by implementing the ALM process in (33);
7:     Optimize fRIS element positions via Algorithm 1;
8:  end while
9:  return  Optimized variables $\mathbf{W}$, $\boldsymbol{\Theta}$, $\mathbf{p}$, $\omega$, $\beta$ and the objective function value.
IV Simulation Results
This section provides numerical results to validate the effectiveness of the fRIS-aided ISAC system and the proposed algorithm. Unless stated otherwise, the simulation parameters are set as follows. The center frequency of this system is $2.4$GHz, corresponding to a wavelength of $\lambda\!=\!0.125$m. The ISAC BS is located at ($3$m, 00m, 00m), and equipped with $M\!=\!8$ antennas. The transmitter has a maximum transmit power of $10$dBm. The fRIS, consisting of $16$ elements ($N_{x}\!=\!N_{y}=4$), is positioned at (00m, $3$m, $3$m), with a size of $A\!\times\!A$ where $A\!=\!4\lambda$ and a minimal inter-element spacing of $\Delta D=\lambda/2$. Additionally, we consider $K\!\!=\!\!4$ communication users and $T\!\!=\!\!3$ targets. The users are randomly distributed within an area centered at ($30$m, $100$m, 00m) with a radius of $10$m. As observed from fRIS, the targets’ (azimuth, elevate) angles are ($-60^{\circ}$, $0^{\circ}$), ($10^{\circ}$, $0^{\circ}$) and ($55^{\circ}$, $0^{\circ}$), respectively. The noise power of the users are set as $\sigma_{0}^{2}\!=\!-60$dBm and QPSK is used for modulation. The weighting coefficient $\alpha$ for balancing the communication and sensing performance is $0.5$. Moreover, the channel fading coefficients $\zeta_{G}$ and $\zeta_{k}$ are calculated by $\eta\cdot(\frac{1}{\text{dist}})^{2}$, where $\eta\!=\!-10$dB is the path loss exponent and dist denotes the transmitter-receiver distance [39]. The penalty factor for ALM and the convergence threshold are set as $1$ and $10^{-5}$, respectively. All simulations are averaged over $100$ Monte Carlo trials.
Figure 2: Convergence behaviors of the proposed algorithm and the benchmarks: the objective function value $\varepsilon$ versus the iteration step.
Figure 3: The effect of RIS element number: the objective function value $\varepsilon$ versus the element number.
The performance of the proposed design is evaluated against the following benchmark schemes:
  • Conventional RIS-aided ISAC system (Conven): The positions of RIS elements are fixed and uniformly distributed across the fRIS plane. The rest of the beamforming designs follow the same procedures as proposed.
  • Discrete position selection (DPS): The moving region of the elements is quantized into discrete grids of equal spacing, where greedy algorithm is employed for the position selection. Specifically, the position of each element is selected sequentially in the discrete position set, while the positions of the other elements are kept unchanged.
  • Random RIS element deployment (Rand): The positions of RIS elements remain fixed for each channel realization but randomly vary across realizations. The optimization of other components is performed as the proposed scheme.
Figure 4: Beampattern at the fRIS: beam gain (dB) versus spatial angle (degree). (a) The beampattern when bias to the communication ($\alpha=0.1$). (b) The beampattern when no bias ($\alpha=0.5$). (c) The beampattern when bias to the sensing ($\alpha=0.9$). (d) The beampattern of different approaches when $\alpha=0.5$.
TABLE I: Comparisons in terms of Average Performance Value
and Time Complexity per round
Scheme Objective Function Value Elapsed Time (s)
Proposed 0.378 313.52
($N=25$)
Conven 0.394 934.01
($N=64$)
Conven 0.432 372.46
($N=49$)
In Fig.2, we examine the convergence performance of our proposed algorithm for the fRIS-assisted ISAC system and compare it with the benchmark scenarios. As expected, the proposed approach demonstrates convergence, with the objective function value decreasing rapidly and stabilizing at $0.57$ after $15$ iterations. Meanwhile, it can be observed that the joint MSE of the proposed fRIS-aided ISAC case outperforms the baselines, especially compared to the conventional RIS-aided ISAC case. This improvement can be attributed to the continuous and flexible element position optimization, which provides additional DoFs in the spatial domain.
Fig.3 illustrates the relationship between the number of fRIS elements ($N$) and the objective function value ($\varepsilon$), with $N$ varying from $16$ to $64$. As observed, the objective function value of the proposed scheme and the baselines all decrease as $N$ increases, while the proposed scheme achieves the best performance among all. In addition, in Table I, we compare the average objective function value $\varepsilon$ and the time complexity of the proposed scheme with those of the conventional RIS-aided ISAC system. Specifically, we select the proposed scheme with $N=25$, the “Conven” scheme with $N=49$ and $N=64$, where either the objective function value or the elapsed time has similar performance. When compared with the “Conven” scheme with $N=64$, the proposed scheme with $N=25$ can achieve slightly better objective function value yet requiring $1/3$ of the elapsed time. When compared with the “Conven” scheme with $N=49$, the proposed scheme with $N=25$ exhibits comparable elapsed time but significantly better objective function value performance. These results demonstrate that employing the fRIS not only compensates for performance losses caused by reducing RIS elements but also alleviates the complexity of high-dimensional optimization.
To evaluate the system’s sensing performance, we further investigate the fRIS beampattern under different weighting coefficients, and compare with the ideal case and the baselines in Fig.4. It is worth noting that the targets are assumed to lie at an angle of $0^{\circ}$ in elevate, so that the beampatterns of the cross section at $0^{\circ}$ in elevate are plotted for clearer observations and easier comparisons. In Fig.4 - Fig.4, the practical, sensing-only and ideal beampatterns are illustrated for the weighted coefficients $\alpha=0.1,0.5\text{ and }0.9$, respectively. It can be observed that beams are consistently formed in the target directions across all $\alpha$ values. However, in the case of $\alpha=0.1$ where the system is biased towards communication, the beam is not as similar as the sensing-only case and exhibits higher sidelobe levels. In contrast, when $\alpha=0.5$ and $\alpha=0.9$ where the system is increasingly biased towards the sensing, the practical beams gradually align with the sensing-only case. On the other hand, Fig.4 compares the beampattern obtained from different approaches when $\alpha=0.5$. It can be found that all the beampatterns have formed peaks in the target directions, but the proposed method achieves lower sidelobe levels compared to the conventional case and higher peaks compared to the ”DPS” and ”Rand” cases.
Figure 5: Average BER (dB) versus different receive noise power (dBm) under different modulations and weighting coefficients.
Figure 6: Average BER (dB) versus different receive noise power (dBm) with different approaches.
In addition, we evaluate the communication performance of the system under different settings. Fig.5 depicts the average bit error rate (BER) of the proposed scheme versus various noise power levels, ranging from $-40$dBm to $-80$dBm. As expected, the average BER decreases when the noise power at the receivers decreases, and QPSK modulation exhibits lower BER when compared to 16QAM. Moreover, when $\alpha$ decreases, the BER further improves as the system becomes more communication-biased. Notably, at a noise power of $-80$dBm and a weighting coefficient of $\alpha=0.9$, the BER can drop to approximately $-28$dB with QPSK modulation. On the other hand, Fig. 6 compares the average BER of the proposed method to the benchmarks. As shown, all cases exhibit a decreasing BER trend as noise power level reduces, while the proposed method consistently achieves a lower BER compared to the baselines.
Figure 7: The effect of the fRIS region size: the objective value versus the normalized region size $A/\lambda$.
In Fig.7, the impact of the normalized region size $A/\lambda$ on the proposed scheme is illustrated, where $A$ varies from $2\lambda$ to $6\lambda$. The results show that the objective function value $\varepsilon$ decreases as the region size increases, which is expected since the larger size of the fRIS provides more DoFs for the mobility of the elements. On the other hand, the benchmarks of “DPS” and “Rand” cases demonstrate a decreasing trend with increasing region size, while the “Conven” case has little change as the region size varies. Compared with the benchmarks, the proposed scheme and the optimization algorithm achieve superior performance across all region sizes, which is consistent to the previous simulation results.
V Conclusions
This paper proposed a fRIS-aided ISAC system to achieve multi-user communications and multi-target sensing by exploiting the beamforming and the fRIS element position optimization. We formulated a joint communication symbol estimation MSE and sensing beampattern mismatch minimization problem subject to transmit power budget, constant-modulus constraint and the element position requirements of the fRIS. An AM-based algorithm was developed, where each subproblem utilized the techniques of derivatives, SDR, ALM and MM optimization. Simulation results validated that the fRIS can significantly reduce the sensing beampattern mismatch and communication symbol estimation error in the ISAC systems, as compared to the conventional RIS.
Appendix A Derivation of the function $\widetilde{f}_{0}$ in (III-E)
To simplify the objective function (34a), we first expand the two square norms, and it is given as
$\!\!\!\alpha\cdot\left(\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}-2\mathfrak{R}\left\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}\right)\!+\!(1\!-\!\alpha)\cdot$
$\!\!\!\!\left(\omega^{2}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\!\!-\!\!2\omega\mathfrak{R}\left\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}\right).$ (49)
It is noteworthy that $\mathbf{s}_{r}^{\mathrm{H}}\mathbf{s}_{r}$ and $\mathbf{s}_{c}^{\mathrm{H}}\mathbf{s}_{c}$ are constants, which have been omitted.
Observing the first term $\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$, it is actually unrelated to the variable $\mathbf{p}$ due to the following reason. Given the property of $\boldsymbol{\Theta}$ that $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ and the expression of $\mathbf{G}$ in (10), this term can be written as $\zeta_{G}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$. Since $\mathbf{a}_{r}(\mathbf{p})$ has the structure of array manifold of (8), $\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})=N$ always holds regardless of element positions. As a result, this term eventually becomes $\zeta_{G}N\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$, which is unrelated to $\mathbf{p}$ and ignorable.
For the second term $-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}$, the expression of $\mathbf{G}$ can also be substituted, yielding $-2\sqrt{\zeta_{G}}\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\}$. Leveraging the cyclic property of trace, it can be re-expressed as $-2\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, where $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$.
Focusing on the third term, it can be rewritten as
$\omega^{2}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$ (50a)
$\overset{\mathrm{(a)}}{\longrightarrow}$ $c_{0}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}\mathbf{A}_{rc}(\mathbf{p})\boldsymbol{\Sigma}_{rc}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})$ (50b)
$\overset{\mathrm{(b)}}{\longrightarrow}$ $c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}\mathbf{a}_{c,k}(\mathbf{p})\mathbf{a}_{c,k}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})$ (50c)
$\overset{\mathrm{(c)}}{\longrightarrow}$ $c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p}).$ (50d)
In operation (a), the expression of $\mathbf{G}$ in (10) and $\mathbf{H}_{rc}$ in (11) are substituted, while $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$ and $c_{0}=\omega^{2}\zeta_{G}c_{x}$ incorporate the variable-independent terms for simplicity. In operation (b), it is derived based on the matrix structure in (14a) and (14b). The operation (c) leverages the diagonal property of $\boldsymbol{\Theta}$, where we denote $\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{c,k}(\mathbf{p}))$.
As for the last term in (A), it can be transformed as follows:
$-2\omega\mathfrak{R}\left\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}$ (51a)
$\overset{\mathrm{(a)}}{\longrightarrow}$ $-2\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\right\}$ (51b)
$\overset{\mathrm{(b)}}{\longrightarrow}$ $-2\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathrm{tr}\left(\mathbf{S}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{r}(\mathbf{p})\right)\right\},$ (51c)
where (a) is the substitution operation, while (b) utilizes the cyclic property of trace and the diagonal property of $\boldsymbol{\Theta}$. We also denote $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
By integrating all the simplified terms, the objective function can be expressed as (III-E). This ends the derivation.
Appendix B Derivation of the function $\widetilde{f}_{1}$ in (III-E)
For the function (III-E), it can be further simplified as follows if one considers the position of $n$-th element position $\mathbf{p}_{n}$ while fixing others. First, examining the first term $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, it is the sum of $N$ number, which can be expressed as
$\sum_{n=1}^{N}-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}e^{\jmath\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})}\right\}$ (52a)
$\overset{\mathrm{(a)}}{\longrightarrow}$ $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}e^{\jmath\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})}\right\}$ (52b)
[…]
In particular, $[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}$ is $n$-th entry of $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\rho_{n}=|[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}|$ and $\xi_{n}=\angle([\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n})$. Additionally, $\nu_{n}=2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is a variable-independent value. The symbol $d_{r,n}(\mathbf{p}_{n})$ is determined by $\mathbf{p}_{n}$, which is given as $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})+p_{y,n}\sin(\psi_{r})$. Moreover, in the operation (a), only the term related to $\mathbf{p}_{n}$ is kept since other element positions are considered constant.
For the second term, one can rewrite it in element-wise manner, i.e.,
$(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})$ (53a)
$\longrightarrow$ $(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\sum_{i=1}^{N}\sum_{j=1}^{N}\zeta_{k}\left[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\right]_{i,j}\!\!\!\cdot\left[\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\right]_{i}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\cdot\left[\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\right]_{j}$ (53b)
$\longrightarrow$ $(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\sum_{i=1}^{N}\sum_{j=1}^{N}\zeta_{k}\left[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\right]_{i,j}\!\!\!$
$\!\!\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}\!\cdot\!e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,j}(\mathbf{p}_{j})-d_{c,k,j}(\mathbf{p}_{j})\right)},$ (53c)
where $d_{c,k,n}(\mathbf{p}_{n})=p_{x,n}\sin(\phi_{c,k})\cos(\psi_{c,k})+p_{y,n}\sin(\psi_{c,k})$. Apparently, when $i=j$, the complex exponential terms are mutually canceled out and become constants unrelated to $\mathbf{p}$. On the other hand, it can be found that the terms of $i\neq j$ appear in conjugate pairs since $\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}$ is a Hermitian matrix, Owing to this characteristics, when considering $n$-th element (i.e. $i=n\text{ or }j=n$) and ignoring the constant terms of $i=j$, (53c) can be reduced to
$(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\zeta_{k}\sum_{i=1,i\neq n}^{N}2\mathfrak{R}\left\{\left[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\right]_{i,n}\right.\!\!\!$
$\!\!\left.\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}\cdot e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}.$ (54)
Notably, when fixing the positions of other elements, the term $[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}]_{i,n}\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}$ is unrelated to $\mathbf{p}_{n}$, which can be viewed constant as $\widetilde{c}_{i,n,k}$. We denote the phase of $\widetilde{c}_{i,n,k}$ as $\widetilde{\xi}_{i,n,k}=\angle\widetilde{c}_{i,n,k}$ while $\widetilde{c}_{i,n,k}$ has an amplitude of $1$. As a result, (B) can further be written as
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\!\zeta_{k}\!\!\!\!\sum_{i=1,i\neq n}^{N}\!\!\!\!2\mathfrak{R}\left\{\widetilde{c}_{i,n,k}e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}$ (55a)
$\longrightarrow$ $\sum_{k=1}^{K}\!\sum_{i=1,i\neq n}^{N}\!\!\widetilde{\nu}_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (55b)
where we denote $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})=d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\nu}_{k}=2(1\!-\!\alpha)c_{0}\zeta_{k}$ for simplicity.
Focusing on the last term, it can be transformed as follows:
$-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathrm{tr}\left(\mathbf{S}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{r}(\mathbf{p})\right)\right\}$ (56a)
$\longrightarrow$ $\!\sum_{n=1}^{N}\sum_{k=1}^{K}\!\!-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n}))}\right\}$ (56b)
$\longrightarrow$ $\sum_{k=1}^{K}-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n}))}\right\}$ (56c)
$\longrightarrow$ $\sum_{k=1}^{K}\!-\overline{\nu}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (56d)
where $s_{n,k}=[\mathbf{S}]_{n,k}$ with the amplitude of $\overline{\rho}_{n,k}=|s_{n,k}|$ and the phase of $\overline{\xi}_{n,k}=\angle s_{n,k}$. We also denote $\overline{\nu}_{n,k}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\overline{\rho}_{n,k}$ for brevity.
This completes the derivation.
Appendix C Gradient and Hessian matrix Derivation of $\widetilde{f}_{1}$
Given that $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the sub-functions $\widetilde{f}_{1,1}$, $\widetilde{f}_{1,2}$, $\widetilde{f}_{1,3}$ shown in (III-E), the derivative of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is equal to the respective derivatives of the sub-functions.
For the first sub-function $\widetilde{f}_{1,1}$, the first-order derivative with respect to $p_{x,n}$ and $p_{y,n}$ is the derivative of a composite function, where the outer function is a cosine function while the interior function is linear. As a result, the can be handled by the chain rule of derivative. First, we derive the the first-order derivatives of $\widetilde{f}_{1,1}$:
$\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{x,n}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\kappa_{r,x}\sin\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (57a)
$\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{y,n}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\kappa_{r,y}\sin\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (57b)
where $\kappa_{r,x}=\sin(\phi_{r})\cos(\psi_{r})$ and $\kappa_{r,y}=\sin(\psi_{r})$. Based on (57), the second-order derivatives can be derived as
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}^{2}}\!$ $=\!\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,x}^{2}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (58a)
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}\partial p_{y,n}}\!$ $=\!\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}\partial p_{x,n}}$
$=\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,x}\kappa_{r,y}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (58b)
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}^{2}}\!$ $=\!\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,y}^{2}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right).$ (58c)
For the second sub-function $\widetilde{f}_{1,2}$, the first-order derivatives and the second-order derivatives can also be derived as
$\!\!\frac{\partial\widetilde{f}_{1,2}}{\partial p_{x,n}}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{x,k}\sin\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right),$ (59a)
$\!\frac{\partial\widetilde{f}_{1,2}}{\partial p_{y,n}}\!=\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{y,k}\sin\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right),$ (59b)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}^{2}}\!=\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}^{2}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59c)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}\partial p_{y,n}}\!=\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}\partial p_{x,n}}$
$\!=\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}\Delta\widetilde{\kappa}_{y,k}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59d)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}^{2}}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{y,k}^{2}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59e)
where $\Delta\widetilde{\kappa}_{x,k}=\kappa_{r,x}-\kappa_{c,x,k}$ and $\kappa_{c,x,k}=\sin(\phi_{c,k})\cos(\psi_{c,k})$, while $\Delta\widetilde{\kappa}_{y,k}=\kappa_{r,y}-\kappa_{c,y,k}$ and $\kappa_{c,y,k}=\sin(\psi_{c,k})$.
Similarly, the first-order and second-order derivatives of $\widetilde{f}_{1,3}$ can be given as
$\!\frac{\partial\widetilde{f}_{1,3}}{\partial p_{x,n}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{x,k}\sin\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60a)
$\!\frac{\partial\widetilde{f}_{1,3}}{\partial p_{y,n}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{y,k}\sin\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60b)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}^{2}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}^{2}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60c)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}\partial p_{y,n}}\!=\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}\partial p_{x,n}}$
$=\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}\Delta\widetilde{\kappa}_{y,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60d)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}^{2}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{y,k}^{2}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right).$ (60e)
By substituting the results of the first-order and second-order derivatives to (39) and (40), the corresponding gradient and the Hessian matrix of $\widetilde{f}_{1}$ will be obtained.
This finishes the derivations.